Random variable X has the following probability distribution
X=X -2 -1 0 1
2 3
PAX) 0.1 0.2 2 03 3k
Find
the value of k
(ii) Evaluate P(X2)&P(-2<X2)
Find The Cumulative Distribution Of X
(iv)
Evaluate the Mean of X
1
Expert's answer
2020-12-08T06:57:43-0500
"\begin{matrix}
 x & -2 & -1& 0 & 1 & 2 & 3 \\
 p(x) & 0.1 & k & 0.2 & 2k & 0.3 & 3k
\end{matrix}"

(i)

"\sum_ip(x_i)=1"

"0.1+k+0.2+2k+0.3+3k=1"

"6k=0.4"

"k=\dfrac{1}{15}"

(ii)


"\begin{matrix}
 x & -2 & -1& 0 & 1 & 2 & 3 \\ \\
 p(x) & 0.1 & \dfrac{1}{15} & 0.2 & \dfrac{2}{15} & 0.3 & 0.2
\end{matrix}"

"P(X<2)=1-P(X=2)-P(X=3)"

"=1-0.3-0.2=0.5"

"P(-2<X<2)=P(X=-1)+P(X=0)+P(X=1)"

"=\dfrac{1}{15}+0.2+\dfrac{2}{15}=0.4"

(iii)


"F(x) = \begin{cases}
 0, &\text{for } x<-2 \\
 0.1, &\text{for } -2\leq x<-1 \\
 1/6, &\text{for } -1\leq x<0\\
 11/30, &\text{for } 0\leq x<1\\
 0.5, &\text{for } 1\leq x<2 \\
 0.8, &\text{for } 2\leq x<3 \\
 1, &\text{for } x\geq3
\end{cases}"

(iv)


"E(X)=-2(0.1)+(-1)(\dfrac{1}{15})+0(0.2)+1(\dfrac{2}{15})"

"+2(0.3)+3(0.2)=\dfrac{16}{15}"

"E(X)=\dfrac{16}{15}"